# Compelling ReLU Network Initialization and Training to Leverage Exponential Scaling with Depth

### v3

@@ Abstract
A neural network with ReLU activations may be viewed as a composition of piecewise linear functions. For such networks, the number of distinct linear regions expressed over the input domain has the potential to scale exponentially with depth, but it is not expected to do so when the initial parameters are chosen randomly. This poor scaling can necessitate the use of overly large models to approximate even simple functions. To address this issue, we introduce a novel training strategy: we first reparameterize the network weights in a manner that forces the network to display a number of activation patterns exponential in depth. Training first on our derived parameters provides an initial solution that can later be refined by directly updating the underlying model weights. This approach allows us to learn approximations of convex, one-dimensional functions that are several orders of magnitude more accurate than their randomly initialized counterparts.
@@ 1 Introduction
Beyond complementary advances in areas like hardware, storage, and networking, the success of neural networks is primarily due to their ability to efficiently capture and represent nonlinear functions (Gibou et al., 2019). In a neural network, the goal of an activation function is to introduce nonlinearity between the network’s layers so that the network does not simplify to a single linear function. The rectified linear unit (ReLU) has a unique interpretation in this regard. Since it either deactivates a neuron or acts as an identity, the resulting transformation on each individual input remains linear. However, each possible configuration of active and inactive neurons can produce a unique linear transformation over a particular region of input space. The number of these activation patterns and their corresponding linear regions provides a way to measure the expressivity of a ReLU network, and can theoretically scale exponentially with the depth of the network [^1] The reader is referred to Chmielewski-Anders (2020) for definitions of linear regions, activation patterns, and activation regions.. Hence deep architectures may outperform shallow ones.
@@
Surprisingly, though, a sophisticated theory of how to best encode functions into ReLU networks is lacking, and in practice, adding depth is often observed to help less than one might expect from this exponential intuition. Lacking more advanced theory, practitioners typically use random parameter initialization and gradient descent, the drawbacks of which often lead to extremely inefficient solutions. Hanin and Rolnick (2019) show a rather disappointing bound pertaining to randomly initialized networks: they prove that the average number of linear regions formed upon initialization is entirely independent of the configuration of the neurons, so depth is not properly utilized. They observed that gradient descent has a difficult time creating new activation regions and that their bounds approximately held after training. As we will discuss later, the number of linear regions is not actually a model property that gradient descent can directly optimize. Gradient descent is also prone to redundancy; Frankle and Carbin (2019) show how up to 90% of neurons may ultimately be eliminated from a network without significantly degrading accuracy.
@@
The present work aims to begin eliminating these inefficiencies, starting in a simple one-dimensional setting. Drawing inspiration from existing theoretical ReLU constructions, our novel contributions include a special reparameterization of a ReLU network that forces it to maintain an exponential number of activation patterns over the input domain. We then demonstrate a novel pretraining strategy, which trains these derived parameters before manipulating the underlying matrix weights. This allows the network to discover solutions that are more accurate and unlikely to be found otherwise. Our technique directly rectifies the issues raised by Hanin and Rolnick (2019). Firstly, we initialize immediately with an exponential number of linear regions, which would not be expected to happen otherwise. Secondly, we are not reliant on gradient descent to “discover” new regions; we only need to maintain the existing ones, which is already guaranteed by the reparameterization during the pretraining stage. Additionally, this pretraining step can avoid areas of the loss landscape where unassisted gradient descent might make short-sighted optimizations that eliminate activation regions. Our results demonstrate that minimizing the reliance of network training on unassisted gradient descent can reliably produce error values orders of magnitude lower than a traditionally-trained network of equal size. Although the preliminary results in this theoretical study pertain to one-dimensional convex functions, the paper concludes with our views on extending these theoretical exponential benefits for ReLU networks to arbitrary smooth functions with arbitrary dimensionality, which would have significant practical utility.
@@ 2 Related Work
This work is primarily concerned with a novel training methodology, but it also possesses a significant approximation theoretic component. Our work can be viewed as a first step towards generalizing the approximation to $x^{2}$ we review in this section. The reparameterization we employ modifies this method to become trainable in order to represent other convex one-dimensional functions, and then converts that result back into a matrix representation.
@@ 2.1 Related Work in Approximation
The novel reparameterization strategy contributed by the present work aims to extend this idea, compelling ReLU networks to compose rapidly oscillating triangular signals. The key insight is that instead of expressing each layer by its weights, we can train the location on $(0,1)$ of a triangular function’s peak and then set that layer’s weights accordingly. Composing the layers together will create a variety of dilated triangular waveforms that the network can use in its approximations. We emphasize that while the Telgarsky (2015) approximation uses fixed symmetric triangles and is not trainable, a key insight of our work is to instead leverage non-symmetric triangles, which yields trainable parameters (one per triangle).
Nonetheless, with only symmetric triangle waves, Yarotsky (2017) and Liang and Srikant (2016) demonstrate the ability to construct $y=x^{2}$ on $[0,1]$ with exponential accuracy. To produce their approximation, one begins with $f_{0}(x)=x$, then computes $f_{1}(x)=f_{0}(x)-T(x)/4$, $f_{2}(x)=f_{1}(x)-T(T(x))/16$, $f_{3}(x)=f_{2}(x)-T(T(T(x)))/64$, and so forth, as pictured in Figure 1. As these successive approximations are computed, Figure 1 plots their convergence to $x^{2}$, as well as the convergence of the derivatives. We note that these prior papers explicitly focus on approximating $x^{2}$ using symmetric triangle waves, unlike the present work, which is more general (although those papers leverage their $x^{2}$ approximation to also indirectly build approximations of other functions).
The techniques of Yarotsky (2017) and Liang and Srikant (2016) seem to touch on something fundamental regarding how ReLU networks can approximate functions: even though individual neuron outputs may be jagged (e.g., Figure 1, top row), an appropriate infinite sum of them (e.g., $f_{\infty}(x)$) may still be differentiable. Moreover, the derivative of the approximation produced by this technique also converges to the true solution $y=2x$ in the limit. Therefore, inspired by the exponential accuracy of this technique, we place a particular emphasis on investigating the relative importance of differentiability throughout the paper. We will argue with numerical evidence that enforcing convergence of the derivative can reduce the number of local minima during training.
@@
One of the appealing properties of neural networks is that they are highly modular. The $x^{2}$ approximation is used by other theoretical works as a building block to guarantee exponential convergence rates in more complex systems. One possible use case is to construct a multiplication gate. Perekrestenko et al. (2018) does so via the identity $(x+y)^{2}=x^{2}+y^{2}+2xy$. The squared terms can all be moved to one side, expressing the product as a linear combination of squared terms. They then further assemble these multiplication gates into an interpolating polynomial, which can have an exponentially decreasing error when the interpolation points are chosen to be the Chebyshev nodes. Polynomial interpolation does not scale well into high dimensions, so this and papers with similar approaches will usually come with restrictions that limit function complexity: Wang et al. (2018) requires low input dimension, Montanelli et al. (2020) uses band limiting, and Chen et al. (2019) approximates low dimensional manifolds. In light of these prior papers, the present work revisits the $x^{2}$ approximation, seeking a generalization that can avoid the limitations of these works. Replacing $x^{2}$ with a more flexible class of functions may produce a suitable building block for techniques that can bypass polynomial interpolation.
@@
Lastly, other works focus on showing how ReLU networks can encode and subsequently surpass traditional approximation methods (Lu et al., 2021; Daubechies et al., 2022). Interestingly, certain fundamental themes from above like composition, triangles, or squaring are still present. One other interesting comparison of the present work is to Ivanova and Kubat (1995), which uses decision trees as a means to initialize neural networks. It is a sigmoid/classification analogy to this work, but rather than an attempting to improve neural networks with decision trees, it is an attempt to improve decision trees with neural networks.
@@ 2.2 Related Work in Initialization
Our work seeks to improve network initialization by making use of explicit theoretical constructs. This stands in sharp contrast the current standard approach, which treats neurons homogeneously. Two popular initialization methods implemented in PyTorch are the Kaiming (He et al., 2015) and Xavier initialization (Glorot and Bengio, 2010). They use weight values that are sampled from distributions defined by the input and output dimension of each layer. Aside from sub-optimal approximation power associated with random weights, a common issue is that an entire ReLU network can collapse into outputting a constant value at initialization. This is referred to as the dying ReLU phenomenon (Qi et al., 2024; Nag et al., 2023). It occurs when the initial weights and biases cause every neuron in a particular layer to output a negative value. The ReLU activation then sets the output of that layer to 0, blocking any gradient updates. Worryingly, as depth goes to infinity, the dying ReLU phenomenon becomes increasingly likely (Lu et al., 2020). Several papers propose solutions: Shin and Karniadakis (2020) use a data-dependent initialization, while Singh and Sreejith (2021) introduce an alternate weight distribution called RAAI that can reduce the likelihood of the issue and increase training speed. We observed during our experiments that RAAI greatly reduces, but does not eliminate the likelihood of dying ReLU. Our approach enforces a specific network structure that does not collapse in this manner.
3 Compositional Networks
We begin by discussing how to deliberately architect the weights of a 4-neuron-wide, depth $d$ ReLU network to induce a number of linear segments exponential in $d$. Throughout the paper, we refer to these as compositional networks. First we will introduce their mathematical form, and then we will discuss how to encode them as ReLU networks.
We define a triangle function as
@@ 3 Compositional Networks
where $0\leq a_{i}\leq 1$. This produces a triangular shape with a peak at $x=a_{i}$ and with both endpoints at $y=0$. Each layer in a compositional network in isolation would compute these if directly fed the input signal. $T_{i}(x)$’s derivatives are the piecewise linear functions:
@@
In a compositional network, the layers feed into each other. This composes triangle functions into triangle waves:
@@
The output of a compositional network will be a weighted sum over the triangle waves formed at each layer. Assuming the network to be infinitely deep, we have
@@
Figure 2: The rightmost diagram is a network representation of a triangle function. The middle shows that triangle function as a hidden layer of a compositional network. The one-dimensional input and output of a triangle function is instead converted into shared weights within the layers. A full compositional network is assembled on the left.
To encode these functions into a ReLU network, we begin with triangle functions. This is the subnetwork on the right in Figure 2. Its maximum output is $1$ at the peak location $a\in(0,1)$. Neuron $t_{1}$ simply preserves the input signal. Meanwhile, $t_{2}$ is negatively biased, deactivating it for inputs less than $a$. Subtracting $t_{2}$ from $t_{1}$ changes the slope at the point where $t_{2}$ begins outputting a nonzero signal. The weight $-1/(a-a^{2})=-(1/a+1/(1-a))$ is picked to completely negate $t_{1}$’s positive influence, and then produce a negative slope. When these components are combined into a deep network, the individual triangles they form will be composed with those earlier in the network, but the two neurons will still behave analogously. Considering the output waveform of $t_{1}$ and $t_{2}$ neurons over the entire input domain $[0,1]$ and at arbitrary depth, we have that $t_{1}$ neurons are always active, outputting complete triangle waves. In contrast, $t_{2}$ neurons are deactivated for small inputs, so they output an alternating sequence of triangles and inactive regions, which can be seen in Figure 3.
@@
Naively using the output of one triangle generator (from Figure 2) as input to the next would form a $1\times 2\times 1\times 2\times 1$ shape, but this is unnecessarily deep. We can replicate the one-dimensional function composition in the hidden layers on the left side of Figure 2 by using weight sharing instead. Any outgoing weight from $t_{1}$ or $t_{2}$ is shared; every neuron taking in a triangle wave as input does so by combining $t_{1}$ and $t_{2}$ in the same proportion. In this way, we can avoid having to use the extra intermediate neurons. Notice in Figure 3 how the waves output by $t_{1}$ and $t_{2}$ neurons always have identical slopes in regions where both neurons are active.
There are two other neurons in the compositional network’s layers. The accumulation neuron (marked as “sum” in Figure 2) acts similarly to a skip connection, maintaining a weighted sum of all previous triangle waves through each layer. If this were naively implemented, it would multiply the $t_{1}$ and $t_{2}$ weights by the sum coefficients. Based on the derivations in the appendix, these coefficients will be exponentially decaying, so learning these weights directly may cause conditioning issues. Instead, all weights in each layer are multiplied by the ratios between successive scaling coefficients, so that the outputs of $t_{1}$ and $t_{2}$ neurons decay in amplitude in each layer. A conventional bias will have no connections to prior layers, so it will be unable to adjust to the rescaling neuron outputs without having to learn exponentially small values. Therefore, a fourth neuron is configured to output a constant signal, so the weight other neurons place on this output can replace their bias. The bias neuron will connect to the previous layer’s bias neuron so that the constant signal can scale down gradually with each layer.

### v4

# Compelling ReLU Networks to Exhibit Exponentially Many Linear Regions at Initialization and During Training

@@ Abstract
A neural network with ReLU activations may be viewed as a composition of piecewise linear functions. For such networks, the number of distinct linear regions expressed over the input domain has the potential to scale exponentially with depth, but it is not expected to do so when the initial parameters are chosen randomly. Therefore, randomly initialized models are often unnecessarily large, even when approximating simple functions. To address this issue, we introduce a novel training strategy: we first reparameterize the network weights in a manner that forces the network to exhibit a number of linear regions exponential in depth. Training first on our derived parameters provides an initial solution that can later be refined by directly updating the underlying model weights. This approach allows us to learn approximations of convex, one-dimensional functions that are several orders of magnitude more accurate than their randomly initialized counterparts. We further demonstrate how to extend our approach to multidimensional and non-convex functions, with similar benefits observed.
@@ 1 Introduction
Beyond complementary advances in areas like hardware, storage, and networking, the success of neural networks is primarily due to their ability to efficiently capture and represent nonlinear functions (Gibou et al., 2019). In a neural network, the goal of an activation function is to introduce nonlinearity between the network’s layers so that the network does not simplify to a single linear function. The rectified linear unit (ReLU) has a unique interpretation in this regard. Since it either deactivates a neuron or acts as an identity, the resulting transformation on each individual input remains linear. However, each possible configuration of active and inactive neurons can produce a unique linear transformation over a particular region of input space. The number of these activation patterns and their corresponding linear regions provides a way to measure the expressivity of a ReLU network [^1] The reader is referred to Chmielewski-Anders (2020) for definitions of linear regions, activation patterns, and activation regions. and can theoretically scale exponentially with the depth of the network (Montufar et al., 2014a; Serra et al., 2018). Hence deep architectures may outperform shallow ones.
@@
Surprisingly, though, a sophisticated theory of how to best encode functions into ReLU networks is lacking, and in practice, adding depth is often observed to help less than one might expect from this exponential intuition. Lacking more advanced theory, practitioners typically use random parameter initialization and gradient descent, the drawbacks of which often lead to extremely inefficient solutions. Hanin & Rolnick (2019) show a rather disappointing bound pertaining to randomly initialized networks: they prove that the average number of linear regions formed upon initialization is entirely independent of the configuration of the neurons, so depth is not properly utilized. They observed that gradient descent has a difficult time creating new activation regions and that their bounds approximately held after training. As we will discuss later, the number of linear regions is not actually a model property that gradient descent can directly optimize. Gradient descent is also prone to redundancy; Frankle & Carbin (2019) show how around 95% of weights may ultimately be eliminated from a network without significantly degrading accuracy.
@@
The present work aims to begin eliminating these inefficiencies, starting in a simple one-dimensional setting. Drawing inspiration from existing theoretical ReLU constructions, our novel contributions include a special reparameterization of a ReLU network that forces it to maintain an exponential number of activation patterns over the input domain. We then demonstrate a novel pretraining strategy, which trains these derived parameters before manipulating the underlying matrix weights. This allows the network to discover solutions that are more accurate and unlikely to be found otherwise. Our technique directly rectifies the issues raised by Hanin & Rolnick (2019). Firstly, we initialize immediately with an exponential number of linear regions, which would not be expected to happen otherwise. Secondly, we are not reliant on gradient descent to “discover” new regions; we only need to maintain the existing ones, which is already guaranteed by the reparameterization during the pretraining stage. Additionally, this pretraining step can avoid areas of the loss landscape where unassisted gradient descent might make short-sighted optimizations that eliminate activation regions. Our results demonstrate that minimizing the reliance of network training on unassisted gradient descent can reliably produce error values orders of magnitude lower than a traditionally-trained network of equal size. Although the preliminary results in this theoretical study pertain to relatively low-dimensional functions, the paper concludes with our views on extending these theoretical exponential benefits for ReLU networks to arbitrary smooth functions with arbitrary dimensionality, which would have significant practical utility.
@@ 2 Related Work
This work is primarily concerned with a novel training methodology, but it also possesses a significant approximation theoretic component. In one dimension, where we initially focus, our work can be viewed as a first step towards generalizing the approximation to $x^{2}$ we review in this section. The reparameterization we employ modifies this method to become trainable in order to represent other convex one-dimensional functions, and then converts that result back into a matrix representation.
@@ 2.1 Related Work in Approximation
The same effect can be achieved with non-symmetric triangle waves (Huchette et al., 2023) (or any shape that sufficiently “folds” the input space (Montufar et al., 2014b)). Our reparameterization strategy focuses on non-symmetric triangle waves. The location on $(0,1)$ of a triangular function’s peak acts as an alternative parameterization of the function, instead of its weight matrix representation. Unlike the raw weights of a depth $d$ ReLU network, any values chosen for the triangle peak parameters will result in the creation of $2^{d}$ linear regions. Furthermore, this parameterization is trainable. By using the peak location to set the raw weights of a layer, the gradients can backpropagate through the raw weights to update the triangle peaks, effectively confining the network to a subspace of weights that generates many linear regions.
The dilated triangular waveforms produced in this manner are not particularly useful on their own. Their oscillations quickly become excessively rapid, and their derviatives do not exist everywhere (especially in the infinite depth limit). But these problems can be rectified by taking a sum over the layers of a network. Yarotsky (2017) and Liang & Srikant (2016) demonstrate the ability to construct $y=x^{2}$ on $[0,1]$ with exponential accuracy using symmetric triangle waves. To produce their approximation, one begins with $f_{0}(x)=x$, then computes $f_{1}(x)=f_{0}(x)-T(x)/4$, $f_{2}(x)=f_{1}(x)-T(T(x))/16$, $f_{3}(x)=f_{2}(x)-T(T(T(x)))/64$, and so forth, as pictured in Figure 1. As these successive approximations are computed, Figure 1 plots their convergence to $x^{2}$, as well as the convergence of the derivative to $2x$. Our reparameterization generalizes this approximation to use non-symmetric triangle waves to approximate a wider class of convex, differentiable, one-dimensional functions.
@@
The $x^{2}$ approximation is used by other theoretical works as a building block to guarantee exponential convergence rates in more complex systems. One possible use case is to construct a multiplication gate. Perekrestenko et al. (2018) does so via the identity $(x+y)^{2}=x^{2}+y^{2}+2xy$. The squared terms can all be moved to one side, expressing the product $xy$ as a linear combination of squared terms. They then further assemble these multiplication gates into an interpolating polynomial, which can have an exponentially decreasing error when the interpolation points are chosen to be the Chebyshev nodes. Polynomial interpolation does not scale well into high dimensions, so this and papers with similar approaches will usually come with restrictions that limit function complexity: Wang et al. (2018) requires low input dimension, Montanelli et al. (2020) uses band limiting, and Chen et al. (2019) approximates low-dimensional manifolds. These works all make use of a fixed representation of $x^{2}$. If our networks were substituted in for the $x^{2}$ approximation, these works would provide theoretical guarantees about the capabilities of the resulting model. Even though the approximation rates will not scale well with input dimension, they serve as a bound that can be improved upon. In Section 5 we further elaborate on how to use our networks to represent higher-dimensional or non-convex functions.
@@
Lastly, other works focus on showing how ReLU networks can encode and subsequently surpass traditional approximation methods (Lu et al., 2021; Daubechies et al., 2022). Interestingly, certain fundamental themes from above like composition, triangles, or squaring are still present. Another interesting comparison of the present work is to Ivanova & Kubat (1995), which uses decision trees as a means to initialize neural networks. It is a sigmoid/classification analogy to this work, but rather than an attempting to improve neural networks with decision trees, it is an attempt to improve decision trees with neural networks. Finally, Park et al. (2021) approaches the problem of linear region maximization from an information theory perspective and uses a loss penalty rather than a reparameterization to increase the number of linear regions.
@@ 2.2 Related Work in Initialization
Our work seeks to improve network initialization by making use of explicit theoretical constructs. This stands in sharp contrast the current standard approach, which treats neurons homogeneously. Two popular initialization methods implemented in PyTorch are the Kaiming (He et al., 2015) and Xavier initialization (Glorot & Bengio, 2010). They use weight values that are sampled from distributions defined by the input and output dimension of each layer. Aside from sub-optimal approximation power associated with random weights, a common issue is that an entire ReLU network can collapse into outputting a constant value at initialization. This is referred to as the dying ReLU phenomenon (Qi et al., 2024; Nag et al., 2023). It occurs when the initial weights and biases cause every neuron in a particular layer to output a negative value. The ReLU activation then sets the output of that layer to 0, blocking any gradient updates. Worryingly, as depth goes to infinity, the dying ReLU phenomenon becomes increasingly likely (Lu et al., 2020). Several papers propose solutions: Shin & Karniadakis (2020) use a data-dependent initialization, while Singh & Sreejith (2021) introduce an alternate weight distribution called RAAI that can reduce the likelihood of the issue and increase training speed. We observed during our experiments that RAAI greatly reduces, but does not eliminate the likelihood of dying ReLU. Our approach enforces a specific network structure that does not collapse in this manner.
3 Initialization and Pretraining Construction
We begin by discussing how to deliberately architect the weights of a 4-neuron-wide, depth $d$ ReLU network to induce a number of linear segments exponential in $d$. As alluded to in Section 2.1, our initialization algorithm is simply to associate a nonsymmetric triangle function with each layer, and then to randomize the location of each triangle’s peak in $(0,1)$; any choice produces $2^{d}$ regions. This yields a reparameterization of the ReLU network in terms of the triangle peak locations, $a_{i}$, rather than in terms of the network’s weights. The pretraining step of our algorithm then involves learning and updating these parameters. This is done by setting individual neuron weights in accordance with the procedures described in this section, and then backpropagating the gradients to update the triangle peak parameters. Pseudocode for our entire algorithm is provided in Appendix A.1.
To illustrate how we set weights, we first describe the mathematical functions that arise in our analysis. Triangle functions are defined as
@@ 3 Initialization and Pretraining Construction
where $10<a_{i}<$ 1. This produces a triangular shape with a peak at $x=a_{i}$ and with both endpoints satisfying $y=0$. Each layer considered in isolation would compute these if directly fed the input signal. $T_{i}(x)$’s derivatives are the piecewise linear functions:
@@
In a deep network, the layers feed into each other, composing their respective triangle functions:
@@
These triangle waves will have $2^{i}$ linear regions, doubling with each layer. The output of the network in pretraining will be a weighted sum over the triangle waves formed at each layer. Assuming the network to be infinitely deep, we have
@@
Figure 2: On the right is a network representation of a triangle function. The middle shows that triangle function as a hidden layer of a network. The one-dimensional input and output of a triangle function is converted into shared weights. A full pretraining network is assembled on the left.
To encode these functions into the weights of a ReLU network, we begin with triangle functions (see the right subnetwork in Figure 2). Its maximum output is $1$ at the peak location $a\in(0,1)$. Neuron $t_{1}$ simply preserves the input signal. Meanwhile, $t_{2}$ is negatively biased, deactivating it for inputs less than $a$. Subtracting $t_{2}$ from $t_{1}$ changes the slope at the point where $t_{2}$ begins outputting a nonzero signal. The weight $-1/(a-a^{2})=-(1/a+1/(1-a))$ is picked to completely negate $t_{1}$’s positive influence, and then produce a negative slope. When these components are combined into a deep network, the individual triangles they form will be composed with those earlier in the network, but the two neurons will still behave analogously. Considering the output waveform of $t_{1}$ and $t_{2}$ neurons over the entire input domain $[0,1]$ and at arbitrary depth, we have that $t_{1}$ neurons are always active, outputting complete triangle waves. In contrast, $t_{2}$ neurons are deactivated for small inputs, so they output an alternating sequence of triangles and inactive regions (see Figure 3).
@@
Naively using the output of one triangle generator (from Figure 2) as input to the next would form a $1\times 2\times 1\times 2\times 1$ shape, but this is unnecessarily deep. We can replicate the one-dimensional function composition in the hidden layers on the left side of Figure 2 by using weight sharing instead. Any outgoing weight from $t_{1}$ or $t_{2}$ is shared; every neuron taking in a triangle wave as input does so by combining $t_{1}$ and $t_{2}$ in the same proportion. In this way, we can avoid having to use the extra intermediate neurons. This is evident in Figure 3 where the waves output by $t_{1}$ and $t_{2}$ neurons posses identical slopes in regions where both neurons are active.
There are two other neurons in the pretraining network’s layers, each with a specific role. The accumulation neuron (marked as “sum” in Figure 2) maintains a weighted sum of all previous triangle waves through each layer. This is similar to a residual connection (He et al., 2016), except that no other neurons take the sum as input (its output is convex, and therefore not rapidly oscillating, and detrimental to making more linear regions). If the sum neuron were naively implemented, it would multiply the $t_{1}$ and $t_{2}$ weights by the sum coefficients. Based on Theorem 4, these coefficients will be exponentially decaying, so learning these weights directly may cause conditioning issues. Instead, all weights in each layer are multiplied by the ratios between successive scaling coefficients, so that the outputs of $t_{1}$ and $t_{2}$ neurons decay in amplitude in each layer. A conventional bias will have no connections to prior layers, so it will be unable to adjust to the rescaling neuron outputs without having to learn exponentially small values. Therefore, a fourth neuron is configured to output a constant signal, so the weight other neurons place on this output can replace their bias. The bias neuron will connect to the previous layer’s bias neuron so that the constant signal can scale down gradually with each layer. We handle this scaling explicitly, rather than relying on regularizers like batch norm (Ioffe & Szegedy, 2015) or layer norm (Lei Ba et al., 2016), partially in order to preserve the mathematical transparency of our model, and partially because this solution arises naturally from our derivations.